OpenStudy (anonymous):

Find the prime factorization of 50. 1 x 2 x 5 x 5 2 x 5 x 5 2 x 25 2 x 2 x 5

OpenStudy (anonymous):

Prime factorization: all the factors of the number in prime numbers, basically means the most simplified form. 1x2x5x5 = 50 = wrong 2x5x5 = 50 = wrong 2x25 = not primed 2x2x5 = ?

OpenStudy (goldphenoix):

What do you mean by, 'wrong'?

OpenStudy (anonymous):

is it the first one? @GoldPhenoix

OpenStudy (goldphenoix):

Yes. How? Well, solve the first one. 1 * 2 = 2 2* 5 = 10 10 * 5 = 50

OpenStudy (anonymous):

o so that's how you do that okay, thank you!

OpenStudy (goldphenoix):

To find the prime factorization of any number, you want to make sure that the factors of it is a prime. Next, you want to make sure that the prime factor multiply to that number.